OpenStudy (anonymous):

HELP! PLEASE completing the square Solve 4x² =- 8x + 12

OpenStudy (anonymous):

Could any of you help me?

OpenStudy (anonymous):

@brookester6 could help me?

OpenStudy (anonymous):

set it to 0

OpenStudy (anonymous):

0=4x^2=-8x+12?

OpenStudy (anonymous):

wait, thats if your factoring lol square each side

OpenStudy (anonymous):

I'm not sure how to do that

OpenStudy (anonymous):

X = -3 x = 1

OpenStudy (anonymous):

What are the steps that I would take to get that answer?

OpenStudy (anonymous):

thats factoring..

OpenStudy (anonymous):

\[4x ^{2}=-8x+12\] \[4x ^{2}+8x-12=0\] Divide by 4 on both sides \[x ^{2}+2x-3=0\] \[x ^{2}-x+3x-3=0\] \[x(x-1)+3(x-1)=0\] \[(x-1)(x+3)=0\] So, x=1 or x=-3

OpenStudy (anonymous):

Solve for x over the real numbers: 4 x^2 = 12-8 x Subtract 12-8 x from both sides: 4 x^2+8 x-12 = 0 The left hand side factors into a product with three terms: 4 (x-1) (x+3) = 0 Divide both sides by 4: (x-1) (x+3) = 0 Split into two equations: x-1 = 0 or x+3 = 0 Add 1 to both sides: x = 1 or x+3 = 0 Subtract 3 from both sides: Answer: | | x = 1 or x = -3

OpenStudy (anonymous):

Aweosme! Thank you all so much. I appreciate it. Very helpful.

OpenStudy (anonymous):

@Brookebrookiee no problem.